Let A⃗ =2i^+3j^, B⃗ =2i^−4j^ , and C⃗ =A⃗ +B⃗ .
1
Expert's answer
2018-10-01T13:57:09-0400
C ⃗=A ⃗ +B ⃗=(2(i ) ̂+3j ̂ )+(2(i ) ̂-4j ̂ )=4(i ) ̂-j ̂
The magnitude is
C=√(4^2+(-1)^2 )=√17≈4.12.
The direction is
θ=-tan^(-1)⁡〖1/4〗=-14° from+x-axis.